# Aperiodic sets of three types of convex polygons

### v3

@@ Appendix D.2 Dividing Tile(1,1)11(1,1)( 1 , 1 ) into convex pentagons and convex hexagons
Let $S_{t}(\delta,\epsilon)$ be the tile set of {$CP(X_{14})$, $CP(X_{3})$, $CP(X_{7})$, $CP(X_{9})$, $CP(X_{11})$} for parameters $\delta$ and $\epsilon$. The convex polygons contained in $S_{t}(\delta,\epsilon)$ become convex pentagonal or hexagonal monotiles if $\delta$ and $\epsilon$ satisfy certain conditions, as listed in Table 1 [3, 11, 12, 18]. All cases in in Table 1 contain the convex pentagonal or hexagonal monotiles that can form periodic tilings without using reflected tiles (see Appendix A and Appendix B)121212 The case of $\epsilon=120^{\circ}$ and $105^{\circ}\leq\delta<165^{\circ}$ also contains the convex pentagonal monotiles belonging only to the Type 2 family which require the use of reflected tiles during the tiling generation process. However, the tile set in the case contains the anterior and posterior sides of convex pentagonal monotiles belonging to the Type 2 family. .
Figure. 29: The method to divide Tile(1,1)11(1,1)( 1 , 1 ) such that the restriction on the number of types of congruent diagrams is satisfied.
Figure. 30: Cases of $\delta+\epsilon=180^{\circ}$, $\epsilon=120^{\circ}$ or $\delta=165^{\circ}$.
Figure. 31: Cases of $\delta+\epsilon=210^{\circ}$, $\delta=135^{\circ}$ or $\epsilon=90^{\circ}$.
@@
Figure. 32: How to divide Tile(1,1)11(1,1)( 1 , 1 ) to create an ASP comprising four types of convex pentagons.
Figure. 33: Example of combination focusing on vertices with an internal angle of $135^{\circ}$.
Figure. 34: Non-periodic tiling generated by the ASP comprising four types of convex pentagons.
@@ Appendix D.4 Conditions for treating anterior and posterior sides as the same type
Figure. 35: Periodic tiling with $CP(X_{14})$, $CP(X_{3})$, and reflected $CP(X_{14})$ as shown in Figure 32. The reflected $CP(X_{14})$ are given an asterisk mark “*.”
@@ Appendix D.5 Supplement 1
In addition to Figure 29, there is another method for dividing Tile(1,1)11(1,1)( 1 , 1 ) such that the number of convex polygonal types is less than or equal to three from a viewpoint that does not distinguish between the anterior and posterior sides, as shown in Figure 36. However, the sets of convex polygons created by the divisions in Figure 36 always contain convex pentagonal monotiles. The convex pentagonal monotile belonging to both the Type 2 and Type 4 families with line symmetry can form periodic tiling (see Appendix A). For the division that makes the convex pentagonal monotiles belonging to the Type 13 family in Figure 36, the tile set of the four types of convex polygons contains the anterior and posterior sides of convex pentagonal monotiles belonging to the Type 13 family. Therefore, the tile sets of convex polygons shown in Figure 36 are not ASP.
Figure. 36: How to divide Tile(1,1)11(1,1)( 1 , 1 ) such that the number of convex polygonal types is less than or equal to three from the viewpoint that does not distinguish between the anterior and posterior sides.
@@
Figure. 37: Example of a method for dividing Tile(1,1)11(1,1)( 1 , 1 ) that does not satisfy the restriction on the number of types of congruent diagrams, even if $ASPcvx(4)$ or $ASPcvx(5)$.
Figure. 38: Example of dividing Tile(1,1)11(1,1)( 1 , 1 ) into four or five convex polygons that could be ASP.
@@ Appendix D.6 Supplement 2
Figure. 39: How to divide Tile(1,1)11(1,1)( 1 , 1 ) containing convex polygons that are candidates for $ASPcvx(4,3)$.

### v5

@@ Appendix D.2 Dividing Tile(1,1)11(1,1)( 1 , 1 ) into convex pentagons and convex hexagons
Let $S_{t}(\delta,\epsilon)$ be the tile set of {$CP(X_{14})$, $CP(X_{3})$, $CP(X_{7})$, $CP(X_{9})$, $CP(X_{11})$} for parameters $\delta$ and $\epsilon$. The convex polygons contained in $S_{t}(\delta,\epsilon)$ become convex pentagonal or hexagonal monotiles if $\delta$ and $\epsilon$ satisfy certain conditions, as listed in Table 1 [4, 15, 16, 24]. All cases in in Table 1 contain the convex pentagonal or hexagonal monotiles that can form periodic tilings without using reflected tiles (see Appendix A and Appendix B)131313 The case of $\epsilon=120^{\circ}$ and $105^{\circ}\leq\delta<165^{\circ}$ also contains the convex pentagonal monotiles belonging only to the Type 2 family which require the use of reflected tiles during the tiling generation process. However, the tile set in the case contains the anterior and posterior sides of convex pentagonal monotiles belonging to the Type 2 family. Thus, they can form the representative tiling of the Type 2 family. .
Figure 29: The method to divide Tile(1,1)11(1,1)( 1 , 1 ) such that the restriction on the number of types of congruent diagrams is satisfied.
Figure 30: Cases of $\delta+\epsilon=180^{\circ}$, $\epsilon=120^{\circ}$ or $\delta=165^{\circ}$.
Figure 31: Cases of $\delta+\epsilon=210^{\circ}$, $\delta=135^{\circ}$ or $\epsilon=90^{\circ}$.
@@
Figure 32: How to divide Tile(1,1)11(1,1)( 1 , 1 ) to create an ASP comprising four types of convex pentagons.
Figure 33: Example of combination focusing on vertices with an internal angle of $135^{\circ}$.
Figure 34: Non-periodic tiling generated by the ASP comprising four types of convex pentagons.
@@ Appendix D.4 Conditions for treating anterior and posterior sides as the same type
Figure 35: Periodic tiling with $CP(X_{14})$, $CP(X_{3})$, and reflected $CP(X_{14})$ as shown in Figure 32. The reflected $CP(X_{14})$ are given an asterisk mark “*.”
@@ Appendix D.5 Supplement 1
In addition to Figure 29, there is another method for dividing Tile(1,1)11(1,1)( 1 , 1 ) such that the number of convex polygonal types is less than or equal to three from a viewpoint that does not distinguish between the anterior and posterior sides, as shown in Figure 36. However, the sets of convex polygons created by the divisions in Figure 36 always contain convex pentagonal monotiles. The convex pentagonal monotile belonging to both the Type 2 and Type 4 families with line symmetry can form periodic tiling (see Appendix A). For the division that makes the convex pentagonal monotiles belonging to the Type 13 family in Figure 36, the tile set of the four types of convex polygons contains the anterior and posterior sides of convex pentagonal monotiles belonging to the Type 13 family. (For the convex pentagonal monotiles belonging to the Type 13 family in Figure 36, the tile set of the convex polygons created by the division contains the reflected tile. Thus, the convex pentagons can form the representative tiling of the Type 13 family.) Therefore, the tile sets of convex polygons shown in Figure 36 are not ASP.
Figure 36: How to divide Tile(1,1)11(1,1)( 1 , 1 ) such that the number of convex polygonal types is less than or equal to three from the viewpoint that does not distinguish between the anterior and posterior sides.
@@
Figure 37: Example of a method for dividing Tile(1,1)11(1,1)( 1 , 1 ) that does not satisfy the restriction on the number of types of congruent diagrams, even if $ASPcvx(4)$ or $ASPcvx(5)$.
Figure 38: Example of dividing Tile(1,1)11(1,1)( 1 , 1 ) into four or five convex polygons that could be ASP.
@@ Appendix D.6 Supplement 2
Figure 39: How to divide Tile(1,1)11(1,1)( 1 , 1 ) containing convex polygons that are candidates for $ASPcvx(4,3)$.